OpenStudy (anonymous):

If I wanted to take the integral of 4/x^2-3, would it be 4ln(x^2-3) divided by the derivative of x^2-3 ?

OpenStudy (badhi):

try differentiating your assumed answer and see whether you get the expression that needs to be integrated

OpenStudy (anonymous):

no this is an arctan

OpenStudy (anonymous):

So a trig substitution

OpenStudy (anonymous):

-4root(3)/3arctan(root3/3)+c

OpenStudy (anonymous):

x inside the arctan

OpenStudy (anonymous):

Wouldnt deriv of arctan be \[1/(x^2+1)\]

OpenStudy (anonymous):

\[-\frac{ 4\sqrt{3} }{ 3 }\arctan(\frac{ \sqrt{2} }{ 3 }x)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i proved that using implicit differentiation, or you may use pyth and trig triangle

OpenStudy (anonymous):

let me give you the link ...

OpenStudy (anonymous):

i proved the derivative so you can see the integral ...

OpenStudy (anonymous):

integral of the above can be done by substituting x=sqrt(3)*sec u

OpenStudy (anonymous):

alternately u can factorise the denominator and apply partial fraction

OpenStudy (anonymous):

no its arctanh

OpenStudy (anonymous):

its a parabola trig

OpenStudy (anonymous):

I got it. I would have to use Partial fractions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but you can use the identity of tanh and just substitute

OpenStudy (anonymous):

the required integral is 2/sqrt(3) ln ((x-sqrt(3))/(x+sqrt(3)))+c

OpenStudy (anonymous):

i say its a tanh

OpenStudy (anonymous):

\[\int\limits \frac{ 4 }{ x^2-3}dx\]

OpenStudy (anonymous):

this gives arctanh